# Deep Reinforcement Learning for Real-Time Drone Routing in Post-Disaster Road Assessment Without Domain Knowledge

### v1

@@ 5.4.2 Testing intance generation
This method preserves the solution space, ensuring interchangeability between original and augmented instances. While previous works (Kwon et al. 2020, Luo et al. 2023, Berto et al. 2024) explored additional transformations (e.g., rotations), we prioritize simplicity and omit these for this study. Future work may investigate their potential to further enhance model performance.
@@ 6.1 Experiment setting
Model Parameters. AEDM is trained with an embedding size $d=128$. Its encoder consists of $l=6$ layers with 8-head MHA and FFN hidden size 512. These parameters choices are informed by successful applications for learning-based methods in solving traditional TSP and VRP (Kool et al. 2018, Kwon et al. 2020, Luo et al. 2023, Zhou et al. 2024, Liu et al. 2024a, Berto et al. 2024). This configuration yields about 1.3 million trainable parameters. Training spans 200 epochs (batch size = 64) on 10,000 on-the-fly generated instances. The Adam optimizer uses an initial learning rate of $10^{-4}$, $L_{2}$ regularization ($10^{-6}$ weight decay), and learning rate decay (×0.1 at epochs 190).
Problem Parameters. AEDM is trained on 100-node instances, comprising 50 nodes in set $\mathcal{P}$ (with damage information values, representing 50 assessable road links) and 50 nodes in set $N$ (without damage information, including one depot). Generalization is evaluated on larger instances (200, 400, 600, 800, and 1,000 nodes), each with 50% nodes in $\mathcal{P}$ and 50% in $N$. Damage values for $\mathcal{P}$ are uniformly sampled from $[1,10]$ and normalized by dividing by 10. Training uses 2, 3, or 4 drones. Practically, drones have a 2-hour battery flight time limit $Q$ and a speed of 60 km/h (Zhang et al. 2023). Given the time-critical nature of disaster response, assessment time limits $p_{\max}$ are set to 30, 45, and 60 minutes (corresponding to 30, 45, and 60 km flight distances). Since $Q$ is significantly larger than $p_{\max}$, and drone flight time being less than $p_{\max}$ inherently implies it is less than $Q$, $Q$ is not considered in subsequent experiments, with focus on $p_{\max}$. Road network coordinates are normalized to $[0,1]$, with link lengths scaled accordingly. For simplified data generation, training uses $p_{\max}=2,3,4$ corresponding to 30, 45, and 60 minutes in real-world distances.
Training time. Training is conducted on Google Colab using an NVIDIA A100 GPU. Each epoch requires approximately 1 minute, resulting in a total training duration of around 3.33 hours for 200 epochs.
@@ 6.2 Benchmark
The proposed problem is novel, with no existing exact or heuristic algorithms tailored to it. For comparison, we developed a mathematical model (detailed in Appendix B. Mathematical formulation) and solved it using commercial solvers (e.g., Gurobi)—a method, like AEDM, that requires no domain knowledge for algorithm design. However, formulating a fully corresponding model for this new problem is complex, demanding domain expertise and significant time investment (see Appendix B. Mathematical formulation).
Given the time sensitivity of disaster response, both solution quality and speed are critical, and we compare both metrics. Notably, disaster scenarios require assessing road networks across multiple locations (e.g., post-earthquake or flood), meaning problem instances to solve are not singular but multiple. Thus, both AEDM and Gurobi must handle multiple instances simultaneously. AEDM leverages GPU parallelism to process up to 1,000 instances per run—consistent with prior applications in combinatorial optimization (e.g., TSP and VRP) (Kool et al. 2018, Kwon et al. 2020, Luo et al. 2023, Zhou et al. 2024, Liu et al. 2024a, Berto et al. 2024). While Gurobi supports parallelization across CPU cores (Zhou et al. 2024, Berto et al. 2024), our experiments—consistent with common real-world deployment constraints due to CPU setup limitations—did not utilize this parallelization for Gurobi. To ensure fairness and reliability despite this constraint, all results are averaged over 10 problem instances for both AEDM and Gurobi. Specifically, AEDM’s total computation time across the 10 instances is reported, while Gurobi’s average time per instance is used—placing Gurobi in an advantageous position, though this does not affect AEDM’s performance.
@@ 6.3 Computational performance
Method #Drone Time (s) Value Gap #Drone Time (s) Value Gap
Gurobi 2 100 11.90 24.23% 3 100 12.73 40.78%
500 13.62 13.29% 500 16.49 23.31%
1,000 14.57 7.26% 1,000 19.10 11.13%
2,000 15.31 2.54% 2,000 21.04 2.14%
AEDM 1 15.07 4.04% 1 20.60 4.18%
$\text{AEDM}\times 8$ 1 15.71 0.00% 1 21.50 0.00%
Gurobi 4 100 15.59 40.62% 5 100 15.62 51.80%
500 20.08 23.50% 500 22.80 29.62%
1,000 20.08 23.50% 1,000 22.80 29.62%
2,000 20.78 20.84% 2,000 24.32 24.92%
AEDM 1 25.12 4.30% 1 31.28 3.45%
$\text{AEDM}\times 8$ 1 26.25 0.00% 1 32.39 0.00%
Table 3: Performance Comparison of Methods in 200-Node Network $p_{\max}=45$ min, Varying Drones)
Method #Drone Time (s) Value Gap #Drone Time (s) Value Gap
Gurobi 2 100 13.40 44.39% 3 100 21.55 33.61%
500 22.02 8.58% 500 26.82 17.37%
1,000 22.05 8.45% 1,000 26.82 17.37%
2,000 22.33 7.29% 2,000 27.82 14.28%
AEDM 1 23.34 3.12% 1 31.33 3.49%
$\text{AEDM}\times 8$ 1 24.09 0.00% 1 32.46 0.00%
Gurobi 4 100 27.62 30.32% 5 100 23.80 47.00%
500 31.97 19.34% 500 34.36 23.47%
1,000 31.97 19.34% 1,000 34.56 23.03%
2,000 31.97 19.34% 2,000 34.76 22.58%
AEDM 1 38.30 3.39% 1 43.29 3.58%
$\text{AEDM}\times 8$ 1 39.64 0.00% 1 44.90 0.00%
@@
Method #Drone Time (s) Value Gap #Drone Time (s) Value Gap
Gurobi 2 100 - 100.00% 3 100 - 100.00%
500 10.22 59.59% 500 11.14 66.63%
1,000 18.61 26.44% 1,000 21.67 35.11%
2,000 22.20 12.26% 2,000 27.79 16.76%
AEDM 1 24.30 3.95% 1 32.06 3.98%
$\text{AEDM}\times 8$ 1 25.30 0.00% 1 33.39 0.00%
Gurobi 4 100 - 100.00% 5 100 - 100.00%
500 17.06 58.32% 500 12.43 73.74%
1,000 21.47 47.56% 1,000 29.10 38.53%
2,000 33.85 17.32% 2,000 35.64 24.72%
AEDM 1 39.00 4.73% 1 44.91 5.12%
$\text{AEDM}\times 8$ 1 40.94 0.00% 1 47.34 0.00%
@@
Method #Drone Time (s) Value Gap #Drone Time (s) Value Gap
Gurobi 2 100 - 100.00% 3 100 - 100.00%
500 27.88 31.31% 500 27.65 45.21%
1,000 31.41 22.63% 1,000 37.30 26.06%
2,000 34.83 14.19% 2,000 42.89 14.99%
AEDM 1 38.96 4.03% 1 48.42 4.04%
$\text{AEDM}\times 8$ 1 40.59 0.00% 1 50.45 0.00%
Gurobi 4 100 - 100.00% 5 100 - 100.00%
500 29.19 52.25% 500 27.39 60.49%
1,000 44.08 27.88% 1,000 43.28 37.58%
2,000 48.99 19.85% 2,000 51.94 25.09%
AEDM 2 58.28 4.66% 2 64.31 7.24%
$\text{AEDM}\times 8$ 2 61.13 0.00% 2 69.33 0.00%
Table 2 and Table 3 compare results across varying numbers of drones in a fixed 200-node network, with the maximum allowable time $p_{\max}$ set to 30 and 45 minutes, respectively. Table 4 and Table 5 present analogous comparisons for a fixed 400-node network under the same $p_{\max}$ values, with drone numbers varied. In each table, the first column lists the names of the methods, where $\text{AEDM}\times 8$ denotes the proposed model with 8× instance augmentation during testing. The second column indicates the number of drones; the third column reports solution time, where Gurobi’s values correspond to its maximum allowed solving time—set to 100, 500, 1,000, and 2,000 seconds to reflect disaster response time constraints. The fourth column shows the optimal objective value at the recorded time, and the fifth column presents the gap, defined as $\text{Gap}=(y-y_{\text{other}})/y$, where $y$ is the objective value from $\text{AEDM}\times 8$ and $y_{\text{other}}$ from other methods. The comparative results reveal three critical insights:
  1. a)
    Solution Quality: AEDM Outperforms Gurobi Across Scales
    Across all node numbers, $p_{\max}$ values, and drone numbers, AEDM and $\text{AEDM}\times 8$ yield higher objective values than Gurobi. For 200-node networks (Table 2Table 3), Gurobi’s solutions show gaps of 2.14–47% relative to $\text{AEDM}\times 8$. For larger 400-node networks (Table 4Table 5), Gurobi’s gaps widen to 12.26–100%, with complete failures in early time windows. Figure 8 visually analyzes these performance gaps across parameter combinations via radar and box-plot charts, confirming $\text{AEDM}\times 8$’s consistent superiority across all metrics, followed by AEDM with minor positive gaps, and Gurobi with the largest deficits.
  2. b)
    Computational Efficiency: AEDM Meets Urgent Time Constraints, Gurobi Lags
    AEDM and $\text{AEDM}\times 8$ achieve inference in 1–2 seconds across all instances, aligning with the urgent time constraints of disaster response. In contrast, Gurobi requires significantly longer computation time (100–2,000 seconds) and fails to produce valid solutions within 100 seconds for 400-node networks (marked by “-” in Table 4 and Table 5).
  3. c)
    Instance Augmentation Enhances Performance
    $\text{AEDM}\times 8$ consistently outperforms the base AEDM, with objective value gaps of 3.12–7.24% across all scenarios, demonstrating that instance augmentation effectively boosts performance.
In summary, AEDM and its augmented variant ($\text{AEDM}\times 8$) emerge as superior solutions for post-disaster road damage assessment, offering unparalleled speed and quality compared to traditional solvers like Gurobi, with instance augmentation further enhancing performance.
Figure 8: Performance Comparison: Average Gap (%) by Method Different Parameter Combinations
@@ 6.4 Method robustness
We further analyze the robustness of AEDM by testing it on parameter distributions that were not seen during training. Figure 9 assesses AEDM’s generalization across drone numbers (2–7), where 2–4 were training-seen and 5–7 were unseen. Gurobi operates within a fixed 2,000-second time limit for all cases. In contrast, AEDM completes real-time optimization across all drone numbers: it maintains high efficiency for training-seen ranges (2–4) and retains fast computation with strong objective values for unseen ranges (5–7). This confirms AEDM’s robustness: it generalizes to unseen distributions, sustaining efficiency and solution quality as problem complexity (unseen drone numbers) increases, outperforming Gurobi’s time-constrained, quality-degrading performance.
@@
Figure 10 assesses AEDM’s robustness across varying maximum allowable assessment time ($p_{\max}$), including 30, 45, and 60 minutes (seen during training) as well as 75, 90, and 105 minutes (held out as unseen). The experimental setup fixes nodes at 400 and drones at 4. As shown in the figure, Gurobi hits a fixed 2,000-second time limit across all $p_{\max}$ settings. In contrast, AEDM completes real-time optimization for all $p_{\max}$ values: it maintains efficiency for training-seen ranges (30–60 minutes) and retains fast computation for unseen ranges (75–105 minutes). Meanwhile, the optimality gap between AEDM and Gurobi widens with increasing $p_{\max}$ (from 17.32% at 30 minutes to 34.92% at 105 minutes). This confirms AEDM’s robustness: it generalizes to unseen $p_{\max}$ distributions, sustaining efficiency and solution quality as temporal complexity increases and outperforming Gurobi’s time-bound, quality-degrading behavior.
Figure 11 assesses AEDM’s robustness on unseen node scales (200–1,000 nodes), with drones fixed at 4 and $p_{\max}=45$. As shown in the figure, for smaller, training-relevant scales (200–400 nodes), AEDM completes real-time optimization. For large, unseen scales (600–1,000 nodes), Gurobi fails to return feasible solutions within the 2,000-second limit (100% time-out gaps, gray bars). In contrast, AEDM still finalizes optimization in real-time, sustaining efficiency even as node number increases. This confirms AEDM’s superiority: it generalizes to large, unseen network scales, maintaining computational feasibility and solution quality. Traditional solvers like Gurobi collapse under scale, highlighting AEDM’s critical advantage for large-scale road network assessments.
@@
In summary, this section evaluates AEDM’s robustness across unseen parameter distributions. Tests on drone numbers, assessment time limits, and node scales show that unlike Gurobi—limited by long computation time and deteriorating results—AEDM sustains efficiency and quality, proving its superiority for time-critical, large-scale road network assessment tasks.

### v2

@@ 5.4.2 Testing instance generation
This method preserves the solution space, ensuring interchangeability between original and augmented instances. While previous works (kwon2020pomo; luo2023neural; berto2024routefinder) explored additional transformations (e.g., rotations), we prioritize simplicity and omit these for this study. Future work may investigate their potential to further enhance model performance.
@@ 6.1 Experiment setting
Model Parameters. AEDM is trained with an embedding size $d=128$. Its encoder consists of $l=6$ layers with 8-head MHA and FFN hidden size 512. These parameters choices are informed by successful applications for learning-based methods in solving traditional TSP and VRP (kool2018attention; kwon2020pomo; luo2023neural; zhou2024mvmoe; liu2024multi; berto2024routefinder). This configuration yields about 1.3 million trainable parameters. Training spans 200 epochs (batch size = 64) on 10,000 on-the-fly generated instances. The Adam optimizer uses an initial learning rate of $10^{-4}$, $L_{2}$ regularization ($10^{-6}$ weight decay), and learning rate decay (×0.1 at epochs 190).
Problem Parameters. AEDM is trained on 100-node instances, comprising 50 nodes in set $\mathcal{P}$ (with damage information values, representing 50 assessable road links) and 50 nodes in set $N$ (without damage information, including one depot). Generalization is evaluated on larger instances (200, 400, 600, 800, and 1,000 nodes), each with 50% nodes in $\mathcal{P}$ and 50% in $N$. Damage values for $\mathcal{P}$ are uniformly sampled from $[1,10]$ and normalized by dividing by 10. Training uses 2, 3, or 4 drones. Practically, drones have a 2-hour battery flight time limit $Q$ and a speed of 60 km/h (zhang2023robust). Given the time-critical nature of disaster response, assessment time limits $p_{\max}$ are set to 30, 45, and 60 minutes (corresponding to 30, 45, and 60 km flight distances). Since $Q$ is significantly larger than $p_{\max}$, and drone flight time being less than $p_{\max}$ inherently implies it is less than $Q$, $Q$ is not considered in subsequent experiments, with focus on $p_{\max}$. Road network coordinates are normalized to $[0,1]$, with link lengths scaled accordingly. For simplified data generation, training uses $p_{\max}=2,3,4$ corresponding to 30, 45, and 60 minutes in real-world distances.
Training time. Training is conducted on Google Colab using an NVIDIA A100 GPU. Each epoch requires approximately 1 minute, resulting in a total training duration of around 3.33 hours for 200 epochs. The source code for AEDM is publicly available at https://github.com/PJ-HTU/AEDM-for-Post-disaster-road-assessment.
@@ 6.2 Benchmark
The proposed problem is novel, with no existing exact or heuristic algorithms tailored to it. For comparison, we developed a corresponding mathematical model (detailed in Appendix B) and solved it using commercial solvers (e.g., Gurobi), which is a method, like AEDM, that requires no domain knowledge for algorithm design. However, formulating a fully corresponding model for this new problem is complex, demanding domain expertise and significant time investment. To provide a more comprehensive evaluation, we additionally designed a two-phase heuristic algorithm based on classical OP frameworks (kobeaga2018efficient; yang2025heuragenix), as detailed in Appendix C. This heuristic consists of a greedy construction phase (GC) followed by a local-search improvement phase (LS). The construction phase uses a classic profit-to-time ratio rule to generate initial solutions, while the improvement phase applies relocate, exchange, and remove–insert operators that are carefully adapted to the link-to-node transformed network structure. Unlike the proposed AEDM, however, this heuristic requires hand-crafted domain knowledge, illustrating the typical limitations of traditional algorithmic approaches in newly emerging problem settings.
Given the time sensitivity of disaster response, both solution quality and speed are critical, and we compare both metrics across all methods: AEDM, Gurobi, and the heuristic. Notably, disaster scenarios require assessing road networks across multiple locations (e.g., post-earthquake or flood), meaning problem instances to solve are not singular but multiple. Thus, AEDM, Gurobi, and the heuristic must all handle multiple instances simultaneously. AEDM leverages GPU parallelism to process up to 1,000 instances per run, which is consistent with prior applications in combinatorial optimization (e.g., TSP and VRP) (kool2018attention; kwon2020pomo; luo2023neural; zhou2024mvmoe; liu2024multi; berto2024routefinder). While Gurobi supports parallelization across CPU cores (zhou2024mvmoe; berto2024routefinder), our experiments, which are consistent with common real-world deployment constraints due to CPU setup limitations, did not utilize this parallelization for Gurobi. The heuristic is implemented in Python and executed on the same hardware platform as other methods. To ensure fairness and reliability despite these constraints, all results are averaged over 10 problem instances for AEDM, Gurobi, and the heuristic. Specifically, AEDM’s total computation time across the 10 instances is reported, while Gurobi’s and the heuristic’s average time per instance are used, placing the traditional methods in an advantageous position, though this does not affect AEDM’s performance.
@@ 6.3 Computational performance
Method #Drone Time (s) Value Gap (%) #Drone Time (s) Value Gap (%)
Gurobi 2 100 11.42 27.31 3 100 12.21 43.21
500 13.11 16.55 500 15.83 26.37
1,000 14.03 10.69 1,000 18.26 15.07
2,000 14.78 5.92 2,000 20.36 5.30
GC 1 11.95 23.93 2 14.24 33.77
GC+LS 18 12.83 18.33 38 16.68 22.42
AEDM 1 15.07 4.07 1 20.60 4.19
$\text{AEDM}\times 8$ 1 15.71 0.00 1 21.50 0.00
Gurobi 4 100 14.92 43.16 5 100 14.98 53.75
500 19.13 27.12 500 21.91 32.36
1,000 19.18 26.93 1,000 22.06 31.89
2,000 20.02 23.73 2,000 23.48 27.51
GC 5 17.83 32.08 10 20.36 37.14
GC+LS 105 19.48 25.79 280 21.58 33.38
AEDM 1 25.12 4.30 1 31.28 3.43
$\text{AEDM}\times 8$ 1 26.25 0.00 1 32.39 0.00
Table 3: Performance Comparison of Methods in 200-Node Network ($p_{\max}=45$ min, Varying Drones)
Method #Drone Time (s) Value Gap (%) #Drone Time (s) Value Gap (%)
Gurobi 2 100 12.90 46.45 3 100 20.76 36.04
500 21.36 11.33 500 25.93 20.12
1,000 21.40 11.17 1,000 25.93 20.12
2,000 21.63 10.21 2,000 26.91 17.10
GC 2 16.31 32.29 3 22.68 30.13
GC+LS 28 19.87 17.52 60 25.76 20.64
AEDM 1 23.34 3.11 1 31.33 3.48
$\text{AEDM}\times 8$ 1 24.09 0.00 1 32.46 0.00
Gurobi 4 100 26.41 33.38 5 100 22.71 49.42
500 30.81 22.28 500 33.10 26.28
1,000 30.88 22.10 1,000 33.28 25.88
2,000 30.95 21.92 2,000 33.45 25.50
GC 7 28.02 29.31 14 30.94 31.09
GC+LS 145 31.60 20.28 385 34.91 22.25
AEDM 1 38.30 3.38 1 43.29 3.59
$\text{AEDM}\times 8$ 1 39.64 0.00 1 44.90 0.00
@@
Method #Drone Time (s) Value Gap (%) #Drone Time (s) Value Gap (%)
Gurobi 2 100 100 3 100 100
500 9.41 62.81 500 10.34 69.03
1,000 17.28 31.70 1,000 20.19 39.53
2,000 20.62 18.50 2,000 25.82 22.67
GC 2 15.98 36.84 4 20.71 37.98
GC+LS 38 18.74 25.93 80 25.09 24.86
AEDM 1 24.30 3.95 1 32.06 3.98
$\text{AEDM}\times 8$ 1 25.30 0.00 1 33.39 0.00
Gurobi 4 100 100 5 100 100
500 15.72 61.60 500 11.41 75.90
1,000 19.96 51.25 1,000 27.17 42.61
2,000 31.84 22.23 2,000 33.47 29.30
GC 10 19.12 53.29 20 26.62 43.77
GC+LS 220 30.08 26.53 520 35.08 25.90
AEDM 1 39.00 4.74 1 44.91 5.13
$\text{AEDM}\times 8$ 1 40.94 0.00 1 47.34 0.00
@@
Method #Drone Time (s) Value Gap (%) #Drone Time (s) Value Gap (%)
Gurobi 2 100 100 3 100 100
500 26.12 35.65 500 25.93 48.60
1,000 29.47 27.40 1,000 35.06 30.51
2,000 32.96 18.80 2,000 41.07 18.59
GC 3 25.82 36.39 6 34.82 30.98
GC+LS 58 33.74 16.88 125 40.48 19.77
AEDM 1 38.96 4.02 1 48.42 4.02
$\text{AEDM}\times 8$ 1 40.59 0.00 1 50.45 0.00
Gurobi 4 100 100 5 100 100
500 27.41 55.16 500 25.61 63.06
1,000 41.13 32.72 1,000 40.31 41.86
2,000 46.22 24.39 2,000 48.33 30.29
GC 14 41.40 32.28 28 47.23 31.87
GC+LS 320 48.88 20.04 840 52.23 24.65
AEDM 2 58.28 4.66 2 64.31 7.24
$\text{AEDM}\times 8$ 2 61.13 0.00 2 69.33 0.00
Figure 8: Performance Comparison Across Methods and Parameter Combinations: (a) Relative Gap Distribution by Parameter Combination (Radar Chart); (b) Average Relative Gap and Computation Time by Method (Box Plot with Logarithmic Time Scale)
[…]
The experimental results demonstrate clear performance advantages of AEDM across all tested scenarios. For 200-node networks, Gurobi exhibits substantial gaps even with extended computation times up to 2,000 seconds, while the heuristic GC+LS achieves improved but still inferior solutions compared to AEDM $\times$ 8. When problem scale increases to 400 nodes, Gurobi fails to return feasible solutions within 100 seconds (marked as “–” in Table 4 and Table 5) and continues to show significant performance gaps even at 2,000 seconds. The heuristic maintains consistent gap percentages but requires computation times that escalate substantially with problem complexity.
In contrast, both AEDM and AEDM $\times$ 8 maintain consistent rapid performance within 1–2 seconds regardless of network size, drone count, or time limit configuration. The augmented version AEDM $\times$ 8 consistently delivers the best objective values through its 8-fold coordinate transformation strategy that explores diverse solution spaces with negligible computational overhead. Figure 8 provides visual confirmation through radar charts and box plots, showing that AEDM $\times$ 8 maintains minimal gaps uniformly across all parameter combinations while other methods exhibit substantially larger deficits. The box plot quantifies average performance across all scenarios, revealing that AEDM achieves only 4.21% average gap with computation time of merely 1.2 seconds, compared to significantly larger gaps and longer times for Gurobi and GC+LS.
An important distinction between AEDM and traditional methods concerns the development effort required. As detailed in Appendix  B and  C, Gurobi and the GC+LS heuristic demanded substantial domain expertise to design problem-specific operators and calibrate search parameters. In contrast, AEDM eliminates such manual algorithm engineering by learning routing policies directly from synthetic training instances. This automated learning paradigm not only achieves superior performance but also provides inherent adaptability when problem constraints evolve.
In summary, these results establish AEDM as a fundamentally superior approach for post-disaster road damage assessment, delivering three critical advantages simultaneously: (1) solution quality that surpasses both exact methods and traditional heuristics by substantial margins (20–71% improvements over Gurobi, 23–35% over GC+LS); (2) rapid computational efficiency with consistent 1–2 second inference regardless of problem scale; and (3) elimination of domain-specific algorithm design requirements through end-to-end learning. This combination positions AEDM as a paradigm shift from expert-driven optimization to data-driven learning for time-critical combinatorial decision-making in humanitarian logistics.
@@ 6.4 Method robustness
We further analyze the robustness of AEDM by testing it on parameter distributions that were not seen during training. Figure 9 evaluates AEDM across varying maximum allowable assessment time ($p_{\max}$), including 45, and 60 minutes that appeared during training as well as 75, 90, and 105 minutes that were not seen before. The experimental setup fixes the number of nodes at 400 and the number of drones at 4. As shown in the figure, Gurobi consistently reaches the 2,000 second time limit for all $p_{\max}$ settings. The two heuristic methods, GC and GC+LS, also exhibit rapidly increasing computational burden as $p_{\max}$ grows. GC provides fast but low quality solutions, while GC+LS improves solution quality with significantly longer runtime. In contrast, AEDM completes rapid computation for all $p_{\max}$ values and remains stable on both training seen ranges (45 to 60 minutes) and unseen ranges (75 to 105 minutes). These results confirm the temporal generalization ability of AEDM and highlight its superiority over traditional algorithms whose performance is hindered by time limits or rapidly increasing complexity.
Figure 10 examines AEDM under varying drone numbers from 4 to 12, with an interval of 2. Drone number 4 is included during training while 6 to 12 are not seen before. The experimental setup fixes the number of nodes at 400 and $p_{\max}=45$. Similar to the previous results, Gurobi reaches the 2,000 second time limit across all settings, and the heuristic methods exhibit instability as drone number increases. GC deteriorates sharply when more drones are required to coordinate, while GC+LS improves on GC but requires computation times that increase from tens to hundreds of seconds. AEDM solves all cases within seconds and maintains strong performance across both seen ranges and unseen ranges. These results demonstrate that AEDM generalizes effectively to higher coordination complexity and sustains both computational stability and solution quality when facing increased combinatorial difficulty.
@@
Figure 11 assesses AEDM on unseen node scales ranging from 200 to 1,000 nodes, with an interval of 200. The experimental setup fixes the number of drones at 4 and $p_{\max}=45$. For smaller scales from 200 to 400 nodes that align with training distributions, GC, GC+LS, and AEDM all produce feasible solutions, although Gurobi reaches the time limit and GC produces low quality results. GC+LS improves the solution quality over GC but with significantly longer computation time. For larger unseen scales from 600 to 1,000 nodes, Gurobi fails to return feasible solutions within the 2,000 second time limit. The heuristic methods also struggle. GC becomes unstable, and GC+LS requires extremely long computation time while still producing lower quality results. AEDM continues to generate solutions within seconds even at 1,000 nodes and produces much higher collected information values than both heuristic baselines. These results demonstrate that AEDM scales effectively to large unseen networks where traditional solvers and heuristics become infeasible.
@@
In summary, this section evaluates AEDM across unseen parameter distributions including assessment time limits, drone numbers, and node scales. Across all settings, AEDM consistently sustains rapid computation and strong solution quality. In comparison, Gurobi and the heuristic methods GC and GC+LS exhibit significant limitations, including long computation time, unstable performance, and failure to return solutions for large scale problems. These findings confirm the strong generalization capability of AEDM and its practical advantage for time critical and large scale road network assessment tasks.